OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. 8, 6, 4, 2, ...

OpenStudy (fibonaccichick666):

so what is happening here?

OpenStudy (fibonaccichick666):

what is the change from 8 to 6?

OpenStudy (anonymous):

These are my answer choices. I've worked the problem and I can't get any of them.

OpenStudy (anonymous):

-2

OpenStudy (fibonaccichick666):

ok, so if we call 8 \(a_0\) that makes six, \(a_1\)

OpenStudy (fibonaccichick666):

so we first have 8-2

OpenStudy (fibonaccichick666):

now how to go from 8 to 4?

OpenStudy (anonymous):

Oh okay, that makes sense. I was making 8 a1 instead of a0.

OpenStudy (fibonaccichick666):

well, the last answer works for 8 being a_1

OpenStudy (fibonaccichick666):

and the first answer works for it being a_0

OpenStudy (anonymous):

So with the first answer to get 4 we would do 8+ -2(3) = 2 so that wouldn't be correct?

OpenStudy (fibonaccichick666):

well with the first answer 4 is a_2

OpenStudy (fibonaccichick666):

with the last, 4 is a_3

OpenStudy (anonymous):

Okay so that would make it 8 + -2(2) = 4 for the first?

OpenStudy (fibonaccichick666):

yea, and the last works if you start with 8=a_1.

OpenStudy (anonymous):

So that would make the last one the answer then? Do you have to start with 8 = a1?

OpenStudy (fibonaccichick666):

no, they start both ways, so it depends on how you define your sequences in class. Do you start with a0 or a1?

OpenStudy (anonymous):

a1 I believe.

OpenStudy (fibonaccichick666):

alright, so, plug in n=1 make sure you get 8, n=2 make sure you get 6 etc. which eq. satisfies this?

OpenStudy (fibonaccichick666):

although I would argue with the teacher if you get it incorrect. They did not specify which to begin with.

OpenStudy (anonymous):

8 + -2(1-1) = 8 8 + -2(1-2) = 6 8 + -2(3-1) = 4

OpenStudy (anonymous):

So would that mean the last one is correct? Also, it's for an online class and it usually isn't always specific which makes things difficult.

OpenStudy (fibonaccichick666):

I would think so. And online classes do suck for this.

OpenStudy (anonymous):

Yeah it does. Thank you for all your help though and for replying so quick!

OpenStudy (fibonaccichick666):

I think most start the sequence with a1, so I'm like 95% confident. np best of luck